OpenStudy (anonymous):

Knowing that -1≤ cos(2x)≤ 1 , how do we know that 0≤ cos^2(2x)≤ 1

OpenStudy (zarkon):

when you square a real number it is non-negative also \[\cos(2x)\le 1\Rightarrow \cos^2(2x)\le1^2=1\] finally it is possible for \(\cos(2x)\) to be zero for example let \(x=\pi/4\) thus \[0\le\cos^2(2x)\le 1\]

OpenStudy (anonymous):

if we square -1≤ cos(2x)≤ 1 , for all sides, the outcome is confusing for me

OpenStudy (anonymous):

sorry to say but i feel sorry for you you have so hard questions i dont understand i wish i could help

OpenStudy (anonymous):

if we square -1≤ cos(2x)≤ 1 , for all sides,what do we get?

OpenStudy (zarkon):

graph each function to see what is going on. ie graph \(\cos(2x)\) and \(\cos^2(2x)\)

OpenStudy (anonymous):

how about algebraically

OpenStudy (anonymous):

You can do it like this: \[ \cos^2 (2x) = \frac {(1+\cos 4x)}{2} \] Now, \(-1 \le \cos 4x \le 1 \implies 0 \le 1+\cos 4x \le 2 \)\[\implies 0 \le \frac {(1+\cos 4x)}{2} \le 1 \implies 0\le \cos 2x \le 1\]

OpenStudy (anonymous):

EDIT: Read the last line as \[ \implies 0 \le \frac {(1+\cos 4x)}{2} \le 1 \implies 0\le \cos^2 (2x) \le 1 \]

OpenStudy (anonymous):

when a cosine function is squared like cos^2x its graph is always +ve and thus its range is from 0<=cos^2x<=1.